Find the generating function of recurrence relation an+1_an=3n ,n less than 0 where ao=1


1
Expert's answer
2021-10-29T02:52:39-0400

Given,

"\implies a_{n+1}-a_n=3n"


characteristic equation:

"1\/x-1\/x^2=0"

"x-1=0"

"x=1"


Homogeneous solution:

"a_h=c\cdot x^n=c\cdot (1)^n=c"


Particular solution:

"a_t=An^2+Bn+C"

"\implies A(n+1)^2+B(n+1)+C-An^2-Bn-C=3n"

"\implies 2An+A+B=3n"

"\implies A=1.5,\ \ B=-1.5"

Hence,

"a_t=1.5n^2-1.5n"


"a_n=a_h+a_t=c+1.5n^2-1.5n"

"a_0=c+1.5(-1)^2-1.5(-1)=1"

"c=-2"


"\boxed{a_n=-2+1.5n^2-1.5n}"